As a result of the transition of an electron between two energy levels light of wavelength 540nm is emitted. Calculate the energy difference between the two levels.
1
Expert's answer
2021-07-18T01:28:01-0400

"E\n\n=\n\nh\n\n\u22c5\n\n\u03bd"

where

E - the energy of the photon

h - Planck's constant, equal to "6.626\u00d710^{\u221234}J s"

ν - the frequency of the photon

Now, notice that you are given the wavelength of the photon, λ

As you know, frequency and wavelength have an inverse relationship described by the equation

"\u03bb\n\n\u22c5\n\n\u03bd\n\n=\n\nc"

, where

c - the speed of light in vacuum, approximately equal to "3\u00d710^8m s^{\u22121}"


"λ

⋅

ν

=

c

⇒

ν

=

\frac{c}



λ"


"E

=

h

⋅

\frac{c}



λ"


This means that the relationship between energy and wavelength looks like this

Another important thing to notice here is that the wavelength of the photon is given in nanometers, nm

. You need to convert this to meters, the unit used for the value of the speed of light.

"E

=

6.626

×

10

^{−

34}

J

s

×

\frac{3

×

10^

8

m

s}



{540

×

10^

{−

9}

m}

=

3.681

×

10^

{−

19}

J"